A spring whose spring constant is 947.7 N/m has been stretched by a force of 5.5 N in the +x direction. Its full stretched length is measured to be 10.6 cm.

What is the original, unstretched length of this spring?
Answer #1

x = F/K

x is change in length

F = 5.5N

K = 947.7N/m

that gives x =0.6 cm

so the original length = 10.6 -0.6 = 10.0 cm
Answer #2

extension,
x = F/k

= 5.5/(947.7)

= 5.8*10^-3 m

= 0.58 cm

L = Lo + x

Lo = L - x

= 10.6-0.58

= 10.02 cm <<<<<------Answer
Answer #3

let the oringinal length of spring be x

then

chage in length = del x = 0.106 - x

Now ,

as Force = K * del x

5.5 = 947.7 *(0.106 - x)

x = 0.1002 m

x = 10.02 cm

unstretched length of this spring is 10.02 cm
Answer #4

By mgh = 1/2kx^2

F*h=1/2kx^2

5.5*0.106=0.5*947.7*x^2

0.583=473.8 x^2

0.0012=x^2

x=0.034 m